OpenStudy (anonymous):

evulate the following double integral [[ 3x-y dy dx where [ is integral sign please help tommowro is exam

OpenStudy (across):

\[\iint(3x-y)dydx.\]First, integrate with respect to y, holding x constant:\[\int(3xy-\frac{1}{2}y^2+c_1)dx.\]Then integrate with respect to x:\[\frac{3}{2}x^2y-\frac{1}{2}xy^2+c_1x+c_2.\]

OpenStudy (anonymous):

you say holding y as constant then y u increase y in first integrating 1/3 y2

OpenStudy (anonymous):

hmm thats good thanku...

OpenStudy (anonymous):

do you treat y as a constant on the the second integration?

OpenStudy (across):

Yes because we're integrating with respect to x.

OpenStudy (across):

Also, altafhpatel, with indefinite integrals it's easy to forget including the constants of integration; you can see above that they're fairly important!

OpenStudy (anonymous):

Thank you ma'am.